how many pounds of salted nuts selling for $4.80 should be mixed with 7 lbs of salted nuts selling for $4.20 per pound to obtain a mixture at $4.38 per pound
Answer by stanbon(75887)

how many pounds of salted nuts selling for $4.80 should be mixed with 7 lbs of salted nuts selling for $4.20 per pound to obtain a mixture at $4.38 per pound
------------------------------------
Equation:
value + value = value
4.8x + 4.2*7 = 4.38(x+7)
Multiply thru by 100 to get rid of the decimals:
480x + 420*7 = 438x + 438*7
42x = 18*7
x = 3 (# of lbs. of $4.80 nuts needed for the mix)
===================================================
Cheers,
Stan H.